OpenStudy (loser66):

how to find second partial derivative of u(x,y) by using u(x+h,y) and u(x,y+h) Please, help

OpenStudy (loser66):

@freckles

OpenStudy (loser66):

\[\dfrac{\partial^2 u}{\partial x^2} +\dfrac{\partial^2u}{\partial y}=\\\dfrac{(x-h,y)-2u(x,y)+u(x+h,y)}{h^2}+\dfrac{(x,y-h)-2u(x,y)+u(x,+hy)}{h^2}\] But I don't know how to get it.

OpenStudy (loser66):

@IrishBoy123

OpenStudy (irishboy123):

well loser, the quotient rule looks to be in there so when i've finished my nightly ablutions, i'm defo up for having a go at this. though i know that @freckles is a much safer bet than i :-) being pedantic with the red.... \[\dfrac{\partial^2 u}{\partial x^2} +\dfrac{\partial^2u}{\partial y}=\\\dfrac{\color{red}{u}(x-h,y)-2u(x,y)+u(x+h,y)}{h^2}+\dfrac{\color{red}{u}(x,y-h)-2u(x,y)+u(x,+hy)}{h^2}\]

OpenStudy (loser66):

Thanks anyway. I am waiting.

OpenStudy (freckles):

\[\lim_{h \rightarrow 0} \frac{f(x+h)-f(x)}{h}=f'(x) \] \[f''(x)=\lim_{h \rightarrow 0} \frac{f'(x+h)-f'(x)}{h} \\ f''(x)=\lim_{h \rightarrow 0} \frac{\lim_{h \rightarrow 0} \frac{f(x+h+h)-f(x+h)}{h}-\lim_{h \rightarrow 0} \frac{f(x+h)-f(x)}{h}}{h}\]

OpenStudy (freckles):

\[f''(x)=\lim_{h \rightarrow 0} \frac{f(x+2h)-f(x+h)-f(x+h)+f(x)}{h^2 } \\ f''(x)=\lim_{h \rightarrow 0} \frac{f(x+2h)-2 f(x+h)+f(x)}{h^2}\] so looks like we need to show this is equivalent to \[f''(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-2f(x)+f(x-h)}{h^2}\] based on what you have above...

OpenStudy (freckles):

I'm thinking you might want to try a sub

OpenStudy (freckles):

yes a sub seems to work

OpenStudy (loser66):

Thanks a lot. I think I got it.